OpenStudy (anonymous):

1)is this rite ? CAN YOU SAY HOW ? =(1-cosA/1+cosA) ^1/2 =1-COSA/SINA

terenzreignz (terenzreignz):

I can see it... you ready? ^^

OpenStudy (anonymous):

ya

terenzreignz (terenzreignz):

Okay, pay attention. \[\Large \left[\frac{1-\cos A}{1+ \cos A}\right]^{\frac12}\]

OpenStudy (anonymous):

oaky

terenzreignz (terenzreignz):

We'll multiply the inside of the bracket by 1, which is ok, except we multiply a rather special variant of "1". Namely \[\large \frac{1-\cos(A)}{1-\cos(A)}\]

terenzreignz (terenzreignz):

So let's do it. \[\Large \left[\frac{1-\cos A}{1+ \cos A}\cdot \frac{1-\cos A}{1-\cos A}\right]^{\frac12}\] Do you see where this is going?

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

Can you take it from here?

OpenStudy (anonymous):

yes i got

terenzreignz (terenzreignz):

So, we're done? ^_^

OpenStudy (anonymous):

but after i want to get =1-cosA/sinA

terenzreignz (terenzreignz):

Okay, relax. Simplify the multiplication, tell me what you get. \[\Large \left[\frac{(1-\cos A)(1-\cos A)}{(1+ \cos A)(1-\cos A)}\right]^{\frac12}\]

OpenStudy (anonymous):

this is my question

terenzreignz (terenzreignz):

I know, but I'm asking you to simplify the multiplication above. ^ You can do it :)

OpenStudy (anonymous):

yes i can do.. after i got finely =[(1−cosA)(1−cosA)(1+cosA)(1−cosA)]12 =[1−cosA1+cosA]12

terenzreignz (terenzreignz):

No, lol... write it down. I can't see clearly when it's all typed out like that

OpenStudy (anonymous):

=[(1−cosA)(1−cosA)(1+cosA)(1−cosA)]12 =[1−cosA1+cosA]12 =1-COSA/sinnA =(1/SINA)-(COSA/SINA) =cosecA*cotA =rhs is this okay my friend?

terenzreignz (terenzreignz):

No, it's like this... one step at a time \[\Large \left[\frac{(1-\cos A)^2}{1- \cos^2 A}\right]^{\frac12}\] Can you now see what's going to happen?

OpenStudy (anonymous):

oh i got this.. [(1-cosa)^2/1-cos^a]^1/2

terenzreignz (terenzreignz):

You forgot the ^2 in the cos in the denominator. Anyway, we can't do anything about the numerator now, but the denominator... you know from Pythagorean identities that that must be equal to...?

OpenStudy (anonymous):

yes now i can write this.. (1-cosA/1+cos^0A )^1/2

terenzreignz (terenzreignz):

Oh... no, you can't. LOL

terenzreignz (terenzreignz):

\[\Large 1-\cos^2A = \color{red}?\] Pythagorean identies... look back.

OpenStudy (anonymous):

sin^2A

terenzreignz (terenzreignz):

That is correct. So denominator now becomes... \[\Large \left[\frac{(1-\cos A)^2}{\sin^2A}\right]^{\frac12}\] Now just get the square root of the whole thing. Finish it! LOL

OpenStudy (anonymous):

terenzreignz oh my god.. thanks you it so ezy now.........lol

terenzreignz (terenzreignz):

No problem :) And call me TJ next time, ok? Signing off now ----------------------------- Terence out